OpenStudy (anonymous):

I bet noone can get this simplify 2+3i _______ 2-5i A.)-11+16 ---------- -21 B.)-11+16i -------- 29 C.)19+16 --------- -21 D.) 1 -3/5i

OpenStudy (anonymous):

\[\frac{2+3i}{2-5i}=\frac{2+3i}{2-5i} .\frac{2+5i}{2+5i}=\frac{-11+16i}{29}=-\frac{11}{29}+\frac{16}{29}i\]

OpenStudy (anonymous):

B

OpenStudy (anonymous):

wow thanks imma give you your 100 medal

OpenStudy (anonymous):

ty :)